OpenStudy (anonymous):

2x +5y =16 and 4x +7y=23

OpenStudy (anonymous):

2x+5y=16 change to 2x=16-5y for first equation, then change 4x+7y=23 to 2x=23/2-7/2y for second equation... then insert first into second 16-5y=23/2-7/2y and equal to 32-10y=23-7y then 3y=9 and y=3 and x=1/2

OpenStudy (anonymous):

Where do you get x has 1/2?

OpenStudy (anonymous):

after get y=3 then you insert y=3 into first equation, then it is 2x=16-5(3) and then 2x=1 and then x= 1/2

OpenStudy (anonymous):

Mariah, my student says thank you very much!

OpenStudy (anonymous):

:)